An object allowed to fall from height of 70m at the same time a projectile is fired straight up at velocity of 35m/s. When and where will they pass each other?
1
Expert's answer
2019-05-17T11:55:36-0400
"Y=70- (g\times t^2)/2"

is the equation of motion of the object


"Y=35\times t-(g \times t^2)/2"

is the equation of motion of the projectile

Equate these equations:


"70- (g\times t^2)/2 =35\times t-(g \times t^2)/2"

From here: t=2sec; Y=50 meter